# Aligning LLMs for FL-free Program Repair

### v1

@@ III D4C: Direct Debug Drives Decent Code
Based on these two insights, we developed an APR framework D4C, i.e., direct debug drives decent code. As shown in Fig. 4, when presented with a new buggy program [^1] In our implementation, the buggy program refer to a single function, due to the limitation of LLMs’ context window length., D4C uses the related artifacts, including the documents and failed test information, to construct a bug report. Then, D4C will use this bug report to instruct the model to generate a refined version of the buggy program. To make LLM follow D4C’s input-output format, we adopt a one-shot prompting strategy by providing a handcrafted input example and its output label. D4C does not require any additional fine-tuning. It only changes the output format to align with the pre-training objectives, and uses test information and documents (e.g., Javadoc comments) in the input prompt to enable repair without the prior knowledge of the correct fault locations. We implemented our approach on GPT-4 series models and Mixtral-MoE models. Next, we will introduce our problem definition (Sec. III-A) and the design details of D4C, including model selection (Sec. III-B), artifact processing (Sec. III-C), prompt construction (Sec. III-D), patch generation (Sec. III-E) and validation (Sec. III-F).
@@ III-A Problem Definition
We model the APR task as a completion task that aims at generating a complete refined program based on a buggy program and associated artifacts. This is different from previous LLM-based APR approaches, which treated the APR task as an infilling or cloze task [14]. Specifically, the optimization objective of D4C can be formally written as:
@@
${\arg\max}_{\theta}\ \ P(t_{\text{Fixed Hunk}}|t_{\text{Buggy Program}};\theta)$ (5)
@@ III-B Model Selection
The LLM itself is always the most vital part of LLM-based applications, which determines the upper bound of the capability for handling specific downstream tasks. Since the goal of proposing D4C is to illustrate an adaptive way to use current advanced decoder-only LLMs for APR, the backbone of D4C should be a decoder-only LLM which has been pre-trained on a substantial code corpus via next token prediction objective. Moreover, D4C is not a complex approach aiming at teaching LLMs to perform APR. Instead, its APR ability primarily comes from aligning the model’s completion format with its pre-training corpus format, e.g., a complete function. Thus, to enable D4C exhibit the state-of-the-art APR performance, we choose the most advanced LLMs (e.g., GPT-4 and Mixtral-MoE) to serve as the backbone in our implementation, as they have long context windows for generating refined version of most of the buggy functions. We will not dive into their architecture or pre-training details in this paper, but it is worth noting that the backbone for D4C can be changed to other LLMs with next token prediction (CLM) training objective. To best ensure our replicability, we employed fixed remote API checkpoints (e.g., gpt-4-0613) or fixed local model versions (e.g., mixtral-moe-instruct-v0.1). For reference, we also provide comparative experiments about using different backbones in Sec. IV-C.
III-C Artifact Processing
@@
Before patch generation, D4C needs to extract related artifacts for subsequent prompting. As shown in Fig. 4, D4C draws (1) program documents or comments that describe the general purpose of the code and its input-output data types, (2) the inputs and expected outputs of the failed test cases, and (3) error messages from those failed test cases, corresponding to the buggy code. Our underlying motivation for providing these artifacts to the LLM is that LLM has gradually exhibited human-like analysis and reasoning abilities recently [29]. Instead of treating LLM as a tool as in the traditional mindset of training NMT models from scratch, we should treat it as an intelligent agent. From this perspective, we contemplate how humans debug, and present this procedure to LLM for execution. Specifically, in the human program repair process, we need to first intuitively understand the function’s purpose so we refer to related code documents or function-level comments. Then, we wonder what input can reproduce the bug. Thus, we will check the relevant failed test cases and error messages. Finally, in the absence of communication or confirmation from other developers (i.e., in a rubber duck debugging [57] cases), we can only use these artifacts to simultaneously locate and repair the bug. Therefore, we extract artifacts and then construct bug report for prompting LLMs to repair buggy program.
@@
In specific deployment scenarios, we extract each buggy program’s test results and documentation through regex matching, and store them in different fields of a table, categorized by the bug ID. It is noteworthy that the implementations of the artifacts processing procedure may slightly differ due to various data characteristics. For instance, although we default to using function-level comments as documents, if certain functions do not have this artifact, D4C can alternatively exploit related README documents. Moreover, if some artifacts are unavailable in some scenarios (e.g., the test cases of the automated testing tools are unknown), D4C will replace these in the table with a placeholder statement, like “This program does not possess any known test cases.
@@ III-D Prompt Construction
Figure 5: The prompt structure of D4C. The details of the code are omitted.
After the extraction of artifacts, D4C assembles these artifacts into a comprehensive bug report. As shown in Fig. 4, the report template is concise, with only the most basic notations of the different components in the bug report (e.g., program and comment), without providing any complex instructions or requirements. This allows LLM to focus more on the bug report, thus accomplishing our goal of inducing and exploiting LLM’s pre-trained knowledge and capabilities for program repair, rather than “teaching” the model to repair. Following this, D4C integrates the bug report into a complete prompt, which is carefully designed to include a concise role-play system instruction (e.g., You are an AI debugger …) and a pre-determined example consisting of a handcrafted bug report and its refined program. Our underlying goal for providing the pre-determined example to the LLM is to stimulate LLM’s in-context learning ability [13], i.e., synthesizing the parallel report-refinement pair in the prompt to further combine the instruction to understand the objective of the APR task and the mapping relation between input and output. This one-shot in-context learning approach also allows us to restrict the output format of the LLM to a refined function instead of generating other output formats related to APR tasks (e.g., unit diff patch [58]). Moreover, in-context learning assists us in automatically extracting the corrected functions from LLM response for automated patch verification. Specifically, as illustrated in Fig. 5, we enhance the output of the example with markdown-style code formatting control symbols (e.g., ‘‘‘java). This enables us to perform regex matching based on these control symbols when extracting the function output for subsequent processes. We do not add more than one example in the prompt, since our goal is to make LLMs follow our designed output format, thereby verifying our insights, rather than using complex prompting strategy to design a powerful APR method. The pre-determined instruction and the example remain fixed and unchanged in the inference stage.
@@ III-E Patch Generation
Unlike traditional APR tools that generate patches specifically for locations that need modification, the patch generation step of D4C refines the entire program in the inference stage. This is primarily because the core goal of D4C is to explore the extent of APR enhancement by aligning the inference objective to decoder-only LLM’s pre-training objective. Furthermore, we do not introduce an additional patch ranking method, which is widely adopted by current APR approaches. This is because D4C only needs to generate at most 10 samples during decoding (i.e., beam search), which is significantly less than the sampling number of existing methods (i.e., over hundreds of sampling). Thus, D4C can subsequently verify all patches instead of the time-consuming process of verifying hundreds of patches. Additionally, although D4C can rank the patches by comparing their naturalness (quantified by perplexity or entropy) [59, 60], a prior study indicates that it is inaccurate to identify the correctness of the program using naturalness [61]. Therefore, we do not use an additional re-ranking strategy to save the computational cost incurred by the perplexity calculation. Similar to most LLM-based techniques that focus on method-level generation [62], we assume that the buggy function is provided where D4C can select lines within the function to modify. In a real-world scenario if the buggy function is not given, we can refine candidate buggy functions by iterating through all functions executed by the failed tests.
@@ III-F Patch Validation
After generating the candidate patches, D4C extracts the refined program and uses them to replace the buggy program in the original source code. Then, D4C runs the corresponding test suite to find plausible patches that compile successfully and pass all tests. Following prior approaches [14, 63, 24], we then manually check the correctness of the plausible patches to find the correct patch (i.e., a patch that is exactly the same or semantically equivalent to the human-written patch).
@@ IV Evaluation
  • RQ1: How does D4C compare against LLM-based APR methods? We compared D4C to state-of-the-art APR methods with perfect FL. We aim to investigate the improvement from aligning output and enhancing input as mentioned in Sec. I.
  • RQ2: How effective are the two insights that drive the design of D4C? We aim to show that modifying the output to a refined function can better align the training objective, and prompting the bug report to locate and repair bugs without FL can enhance APR performance.
@@ IV-A1 Environment and Implementation
We use a black-box LLM (i.e., gpt-4-0613) and a white-box LLM (i.e., Mixtral-8x7B-Instruct-v0.1) via OpenAI APIs [64] and 8xA100 NVIDIA GPU server for program repair. We use Python 3.9 to implement the inference and evaluation scripts in a local machine with Ubuntu 20.04.5 LTS. To enhance LLM’s instruction following ability, we manually crafted a bug-fix example. This example is employed in each experiment to construct 1-shot prompting. To adapt to different experiments, we modified its input-output format according to the corresponding experiment requirement. Following the previous work, we use the default temperature=1.0. However, we adopt a much smaller sampling number of 10, as the inference budgets (fee, time, etc.) for LLMs are much larger than traditional methods. Unless otherwise specified, all experiments adhered to this setting.
IV-A2 Datasets
Our experiments are conducted on 1027 logic single-function bugs from DebugBench (590) and bugs from Defects4J (437). Specifically, DebugBench [35] is a new debug benchmark designed to counter data leakage (by implanting bugs into source data with GPT-4), which contains a total number of 4,253 bugs from Java, C++, and Python3 from LeetCode. As this paper focuses on APR (instead of syntax error repair), we only evaluate the logic bug type in DebugBench. Since DebugBench is not exposed to the threat of data leakage, we use it for our insight validation (RQ2). However, due to the complexity to reproduce current baselines on this new benchmark, we do not use it for the comparison between D4C and baselines. Defects4J [34] is a widely used APR benchmark with 835 real-world bugs from 17 open-source repositories. Following previous work [23, 24, 25], we separate the single-function bugs of Defects4J to v1.2 (203 bugs) and v2.0 (234 bugs). Since it contains data from previous versions of open-source projects where most code models have been trained on, it faces the threat of data leakage. However, this threat is believed to be less severe in APR compared to other domains [17] due to the sparsity of bug-fix parallel corpus in the training data, meaning that model can only learn the individual buggy or fixed version of the program, not their pairs. Despite this, the potential threat of data leakage could still undermine the validity of hypothesis verification. Hence, Defects4J is mainly used as a reference for comparing D4C with existing baselines. The statistic of the benchmarks are shown in Table I.
@@ IV-A3 Metrics
Following the previous work, we use the number of correct patch to evaluate the APR effectiveness. Specifically, if a patch can pass all unit tests, then it will considered as a plausible patch. If this plausible patch truly resolves the bug, rather than overfit to pass the unit test only (e.g., via referring the provided failed tests in the prompt), then it will be confirmed as a correct patch. In our evaluation, we adhere to prior studies [14, 24, 63] and manually identify whether a plausible patch truly resolves the bug. Notably, the bugs from DebugBench are collected from LeetCode, and the LeetCode programs that can pass all the unseen test suites from the website are deemed as correct programs. In our experiment, we can only access the fixed test examples of each bug in our bug report, rather than the failed unseen test from LeetCode. Thus, we also determine the patches that can pass all the LeetCode tests as correct. To distinguish from manually checked correct patches, we refer to patches can pass the LeetCode validation as verified patches. More details of correctness validation are discussed in Sec. V.
@@ IV-A4 Baselines
In RQ1, we selected several state-of-the-art LLM-based APR methods as baselines, including AlphaRepair [14], FitRepair [19], ChatRepair [24], RAP-Gen [18], and Repilot [63]. We did not compare D4C with learning-based APR methods trained from scratch since our goal is to illustrate the effectiveness of D4C among all the LLM-based approaches. All these baselines employ the infilling objective. In RQ2, we chose different input and output format for comparison to show the superiority of aligning output format and providing artifacts in the input. We also chose prompts that add or delete different information (e.g., function comments) as baselines to investigate the contributions of different information to end-to-end APR.
TABLE II: The number of correct patches generated by different approaches on Defect4J.

### v2

@@ III D4C: Direct Debug Drives Decent Code
Based on these two insights, we developed an APR framework D4C, i.e., direct debug drives decent code. As shown in Fig. 4, when presented with a new buggy program [^1] In our implementation, due to the LLM’s token length limitation, a buggy program refer to a buggy function that can be provided by method-level FL., D4C uses the related artifacts, including the documents and failed test information, to construct a bug report. Then, D4C will use this bug report to instruct the model to generate a refined version of the buggy program. To align LLM’s inference objective with its pre-trained completion objective, we adopt a one-shot prompting strategy. This involves prefixing a fixed example of buggy program and its refined version before the target buggy program in the prompt, enabling LLM to infer their input-output relations and finally generate a refined version of target program in expected format. D4C does not require any additional fine-tuning. It only changes the output format to align with the pre-training objectives, and uses test information and documents (e.g., Javadoc comments) in the input prompt to enable repair without the prior knowledge of the correct buggy hunks. We implemented our approach on GPT-4 series models and Mixtral-MoE models. Next, we will introduce our problem definition (Sec. III-A) and the design details of D4C, including model selection (Sec. III-B), artifact extraction (Sec. III-C), prompt construction (Sec. III-D), patch generation (Sec. III-E) and validation (Sec. III-F).
@@ III-A Problem Definition
We model the APR task as a completion task that aims at generating a complete refined program based on a buggy program and associated artifacts. This is different from previous LLM-based APR approaches, which treated the APR task as an infilling or cloze task [14]. Notably, the artifacts are essential for D4C to identify the location to modify, while infilling APR methods rely on statement-level fault localization to pinpoint buggy hunks. Although the artifacts are also used in some approaches [25] to enhance repair, they are not necessarily required by all infilling methods. Specifically, the optimization objective of D4C can be formally written as:
@@
${\arg\max}_{\theta}\ \ P(t_{\text{Fixed Hunk}}|t_{\text{Buggy Program}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t_{\text{Artifacts}}};\theta)$ (5)
@@ III-B Model Selection
The LLM itself is the most vital part of LLM-based applications. Since the goal of proposing D4C is to illustrate an adaptive way to use current advanced decoder-only LLMs for APR, the backbone of D4C should be a decoder-only LLM which has been pre-trained on a substantial code corpus via next token prediction objective. Moreover, D4C does not aim at teaching extra knowledge to LLMs for APR. Instead, its APR ability primarily comes from mining the pre-trained knowledge via aligning the model’s response with its training corpus, e.g., a complete function. Thus, to enable D4C exhibit the state-of-the-art APR performance, we choose the most advanced LLMs, GPT-4, to serve as the backbone in our implementation. However, since GPT-4 is a black-box model whose inference loss of CLM objective (i.e., perplexity) is unavailable, we have to choose another state-of-the-art white-box model, Mixtral-MoE, as alternative backbone to validate whether generating a whole refined function can better align the training objective. We select Mixtral-MoE because at the time of submission, the strongest model on the HumanEval leaderboard [59] with the longest context window was Mixtral-MoE. We will not dive into their architecture or pre-training details in this paper, but it is worth noting that the backbone for D4C can be changed to other LLMs with next token prediction (CLM) training objective. To best ensure our replicability, we employed fixed remote API checkpoints or fixed local model versions. For reference, we also provide comparative experiments about using different backbones in Sec. IV-C.
III-C Artifact Extraction
@@
Before patch generation, D4C needs to extract related artifacts for subsequent prompting. As shown in Fig. 4, once the buggy function is identified, D4C will automated extract (1) documents or comments that describe the general purpose of the function and its input-output data types, (2) the inputs and expected outputs of the failed test cases, and (3) error messages from executing those failed test cases. Our underlying motivation for providing these artifacts to the LLM is that LLM has gradually exhibited human-like analysis and reasoning abilities recently [31]. Instead of treating LLM as a tool as in the traditional mindset of training NMT models from scratch, we should treat it as an intelligent agent. From this perspective, we contemplate how humans debug, and present this procedure to LLM for execution. Specifically, in the human program repair process, we need to first intuitively understand the function’s purpose so we refer to related code documents or function-level comments. Then, we wonder what input can reproduce the bug. Thus, we will check the relevant failed test cases and error messages. Finally, in the absence of communication or confirmation from other developers (i.e., in a rubber duck debugging [60] cases), we can only use these artifacts to simultaneously locate and repair the bug. Therefore, we extract artifacts and construct bug report for LLM to repair.
@@
Notably, the implementations of the artifacts extraction procedure may differ due to various data characteristics. For instance, although we default to using function-level comments as documents, if certain functions do not have this artifact, D4C can alternatively exploit related README documents. Moreover, if some artifacts are unavailable in some scenarios (e.g., the test cases of the online judgment are unknown), D4C will replace these in the table with a placeholder statement, like “This program does not possess any known test cases.
@@ III-D Prompt Construction
Figure 5: The prompt structure of D4C. The details of the code are omitted. The example pair is fixed, which is used to constrain LLM’s response format.
After the extraction of artifacts, D4C assembles these artifacts into a comprehensive bug report. As shown in Fig. 4, the report template is concise, with only the most basic notations of the different components in the bug report (e.g., program and comment), without providing any complex instructions or requirements. This allows LLM to focus more on the bug report, thus accomplishing our goal of inducing and exploiting LLM’s pre-trained knowledge and capabilities for program repair, rather than “teaching” the model to repair. Following this, D4C integrates the bug report into a complete prompt, which is carefully designed to include a concise role-play system instruction (e.g., You are an AI debugger …) and a fixed example consisting of a handcrafted bug report and its refined program. Our underlying goal is to stimulate LLM’s in-context learning (ICL) ability [13], i.e., analogizing the report-refine pair in the prompt to further understand the instruction. This one-shot ICL also allows us to restrict the output format of the LLM to a refined function rather than other common output formats related to APR (e.g., unit diff patch [61]). Moreover, ICL assists us in automatically extracting the corrected functions from LLM response for automated patch verification as illustrated in Fig. 5. We do not add more than one example in the prompt since our goal is to make LLMs follow the output format, thereby verifying the effectiveness of objective alignment, rather than developing a powerful APR method via complex prompting strategy.
@@ III-E Patch Generation
Unlike traditional APR tools that generate patches specifically for locations that need modification, the patch generation step of D4C refines the entire program in the inference stage. This is primarily because the core goal of D4C is to explore the extent of APR enhancement by aligning the inference objective to decoder-only LLM’s pre-training objective. Furthermore, we do not introduce an additional patch ranking method, which is widely adopted by current APR approaches. This is because D4C only needs to generate at most 10 patches during decoding (i.e., beam search), which is significantly less than the sampling number of existing methods (i.e., over hundreds of sampling). Thus, D4C can subsequently verify all 10 patches without involving significant waste of effort on incorrect patches (within 5 minutes), rather than suffering from a time-consuming process of verifying hundreds of patches (up to 5 hours [14, 25, 20, 62, 21]). Additionally, although D4C can rank the patches by comparing their naturalness (quantified by perplexity or entropy) [63, 64], a prior study indicates that it is inaccurate to identify the correctness of the program using naturalness [65]. Therefore, we do not use any re-ranking strategy to save the computational cost incurred by the perplexity calculation. Following previous LLM-based work that focus on method-level generation [66], we assume that the buggy function is provided where D4C can select lines within the function to modify. In a real-world scenario, the buggy function can be provided by existing method-level FL tools [58] via analyzing the execution results of failed tests.
@@ III-F Patch Validation
After generating the candidate patches, D4C extracts the refined program and uses them to replace the buggy program in the original source code. Then, D4C runs the corresponding test suite to find patches that compile successfully and pass all tests. However, due to the potential incomplete test coverage problem of benchmarks, the APR methods may often produce patches that pass all tests without exactly meeting the expected functionality of the developer [67], thus do not truly fix the bug. These patches are known as plausible patches. Ideally, developer validation is helpful, but the bugs in Defects4J are historically fixed and outdated [36], making developer feedback impractical. To ensure rigor experiment, we follow existing APR papers [14, 62, 20, 21, 25] by adding a manual validation for each plausible patch after automated evaluation to identify the correct patches that are the same or semantically equivalent to the human-written patch.
@@ IV Evaluation
  • RQ1: How does D4C compare against LLM-based APR methods? We compared D4C to state-of-the-art APR methods with and without perfect FL. We aim to investigate the improvement from aligning output and enhancing input as mentioned in Sec. I.
  • RQ2: How effective are the two insights that drive the design of D4C? We aim to show that generating a refined function better aligns LLM’s training objective, and using the artifacts to locate and repair bugs without given buggy hunks further exploits LLM’s APR ability.
@@ IV-A1 Environment and Implementation
We use a black-box LLM (i.e., gpt-4-0613) and a white-box LLM (i.e., mixtral-8x7b-instruct-v0.1) via OpenAI APIs [68] and 8xA100 NVIDIA GPU server for program repair. We use Python 3.9 to implement the inference and evaluation scripts in a local machine with Ubuntu 20.04.5 LTS. To enhance LLM’s instruction following ability, we manually crafted a fixed bug-fix example. This example is employed in each experiment to construct 1-shot prompting for all bugs. To adapt to different experiments, we modified its input-output format according to the corresponding experiment requirement. To control a moderated randomness of text generation, we follow the previous work and set the randomness hyper-parameter, temperature, as 1.0. This is also the default value of OpenAI [68]. To facilitate patch validation, we set a timeout threshold of 1 minutes for each patch. Furthermore, we adopt a much smaller sampling number of 10, as the inference budgets (fee, time, etc.) for LLMs are much larger than traditional methods. Unless otherwise specified, all experiments adhered to this setting.
@@
IV-A2 Datasets
Our experiments are conducted on 1027 logic single-function bugs from DebugBench (590) and bugs from Defects4J (437). Specifically, DebugBench [37] is a new debug benchmark designed to counter data leakage (by implanting bugs into source data with GPT-4), which contains a total number of 4,253 bugs from Java, C++, and Python3 from LeetCode. In general, fixing logic bug is more difficult than fixing other types of bugs. For example, syntax bugs can be easily detected by the compiler or interpreter without testing, and their fixing methods are usually provided in the raised exceptions. Thus, our paper aligns with the practice of existing APR methods that only focus on fixing logic bugs. Since DebugBench is not exposed to the threat of data leakage, we use it for our insight validation (RQ2). However, due to the complexity to reproduce current baselines on this new benchmark, we do not use it for the comparison between D4C and baselines. Defects4J [36] is a widely used APR benchmark with 835 real-world bugs from 17 open-source repositories. Following previous work [24, 25, 26], we separate the single-function bugs of Defects4J to v1.2 (203 bugs) and v2.0 (234 bugs). Since it contains data from previous versions of open-source projects where most code models have been trained on, it faces the threat of data leakage. However, this threat is believed to be less severe in APR compared to other domains [17] due to the sparsity of bug-fix parallel corpus in the training data, meaning that model can only learn the individual buggy or fixed version of the program, not their pairs. Despite this, the potential threat of data leakage could still undermine the validity of hypothesis verification. Hence, Defects4J is mainly used as a reference for comparing D4C with existing baselines. The statistic of the benchmarks are shown in Table I.
@@ IV-A3 Metrics
Following the previous work, we use the number of correct patch to evaluate the APR effectiveness. Specifically, if a patch can pass all unit tests, then it will considered as a plausible patch. If this plausible patch truly resolves the bug, rather than overfit to pass the unit test only (e.g., via referring the provided failed tests in the prompt), then it will be confirmed as a correct patch. In our evaluation, we adhere to prior studies [14, 25, 62] and manually identify whether a plausible patch truly resolves the bug. Notably, the bugs from DebugBench are collected from LeetCode, and the LeetCode programs that can pass all the unseen test suites from the website are deemed as correct programs. In our experiment, we can only access the provided test examples of each LeetCode problem in our bug report, rather than the failed unseen test from LeetCode OJ. Thus, we also determine the patches that can pass all the LeetCode tests as correct. To distinguish from manually checked correct patches, we refer to patches can pass the LeetCode validation as verified patches.
@@ IV-A4 Baselines
In RQ1, we selected several state-of-the-art LLM-based APR methods as baselines, including AlphaRepair [14], FitRepair [21], ChatRepair [25], RAP-Gen [20], and Repilot [62]. We did not compare D4C with learning-based APR methods trained from scratch since our goal is to illustrate the effectiveness of D4C among all the LLM-based approaches. All these baselines employ the infilling objective. In RQ2, we chose different input and output format for comparison to show the superiority of aligning output format and providing artifacts in the input. We also chose prompts that add or delete different information (e.g., function comments) as baselines to investigate the contributions of different information to APR.
IV-B RQ1: Comparison against LLM-based APR
We compare D4C with five state-of-the-art LLM-based APR approaches on Defects4J, the most commonly used benchmark for APR. Since some of the baselines are not open-sourced, we cannot re-run them to reproduce the result. Thus, we reuse their Defects4J results reported in the original paper. Although using Defects4J may have the the risk of exposing its data to LLM’s training corpus, we still use it for evaluation because (1) existing work [17] has shown that data leakage is less of a concern for APR compared to other code-related tasks as the training corpus often contains at most the individual versions of the program (i.e., learning buggy or fixed version only), and LLMs can hardly learn the APR tasks without bug-fix pair corpus during training; and (2) the baselines have also been evaluated on Defects4J (hence, they are subject to the same data leakage risk). Following prior work, we retain this setting to ensure fairness of the comparison. The comparison result is shown in Table II and Table III.
TABLE II: The number of correct patches generated by different approaches on Defects4J v1.2 & v2.0.